# Gradient-free training of recurrent neural networks

### v1

@@ 1 Introduction
Recurrent neural networks (RNNs) are notoriously difficult to train because their loss gradients backpropagated in time tend to saturate or diverge during training, commonly referred to as the Exploding and Vanishing Gradient Problem (EVGP) (Pascanu et al., 2013; Schmidt et al., 2019). In Equation 1, we denote a general non-linear system governed by maps $F$ and $g$, controlled by external inputs ${\bm{x}}_{t}$. A typical RNN modeling approach would directly train all parameters of the networks ${\mathcal{F}},{\mathcal{G}}$ as well as all matrices $K,B,C,V$ on the right of Equation 1.
In our approach, we first construct a linear SSM based on Koopman operator theory (cf. Korda and Mezić (2018a)). The matrix $K$ maps states from ${\bm{z}}_{t-1}$ to ${\bm{z}}_{t}$, with the input ${\bm{x}}_{t}$ affecting this dynamic through the matrix $B$. The observations are related to the states through the linear map $V$.
@@
We propose to use a non-linear neural network to map the low-dimensional state ${\bm{h}}$ into a higher dimension, through ${\mathcal{F}}({\bm{h}}):=\sigma(W{\bm{h}}+{\bm{b}})$, and, similarly, to map the input ${\bm{x}}$ to a high-dimensional input ${\mathcal{G}}({\bm{x}})=\sigma_{x}(W_{x}{\bm{x}}+{\bm{b}}_{x})$. This turns the linear SSM into a non-linear recurrent neural network. Generally, such a non-linear map is also the core idea for many Koopman operator-based models: start with the state ${\bm{h}}_{t-1}$ of a non-linear system, use a non-linear map to a higher-dimensional state ${\bm{z}}_{t-1}$, and then move forward in time in this high-dimensional space using a linear map $K$. This map is then an approximation of the linear Koopman operator of the original, non-linear system.
@@
Crucially, in this contribution, we sample the weights $W$ and biases ${\bm{b}}$ from a specific, data-dependent probability distribution. The remaining free parameters in the matrices $K,B,C,V$ can then be found by solving a sequence of linear problems (cf. Figure 1). This approach alleviates all issues related to the backpropagation of gradients and connects the idea of recurrent architectures to the Koopman operator of the underlying dynamical system. The latter allows us to prove convergence results about the approximation quality w.r.t. the number of neurons in ${\mathcal{F}}$. We demonstrate the performance of this approach on several challenging examples with synthetic and real data.
@@ 2 Related work
Exploding and vanishing gradients. For all major types of RNNs, including LSTMs and GRUs, the dynamics and loss gradients of RNNs are closely linked. If the RNN dynamics converge to a stable fixed point or cycle, loss gradients will remain bounded, but they may vanish (Mikhaeil et al., 2022). Yet, established remedies (Hochreiter and Schmidhuber, 1997; Schmidt et al., 2019) can be used to effectively prevent their gradients from vanishing. However, in chaotic dynamics, gradients invariably explode, posing a challenge that cannot be mitigated through RNN architectural adjustments, regularization, or constraints; instead, it necessitates addressing the problem during the training process (Mikhaeil et al., 2022). Bifurcations may also contribute to sudden jumps in loss observed during RNN training, potentially hindering the training process severely (Doya et al., 1992; Eisenmann et al., 2023). Eisenmann et al. (2023) have demonstrated that specific bifurcations in ReLU-based RNNs are always associated with EVGP during training. Therefore, to harness the full potential of RNNs, the training algorithm needs careful design to tackle challenges posed by bifurcations and the possible emergence of EVGP.
Curse of memory. The existence of long-term memory adversely affects the learning process of RNNs (Bengio et al., 1994; Hochreiter et al., 2001; Li et al., 2021). This negative impact is captured by the concept of the “curse of memory”, which states that when long-term memory is present in the data, approximating relationships demands an exponentially large number of neurons, resulting in a significant slowdown in learning dynamics. Specifically, when the target relationship includes long-term memory, both the approximation and optimization of the learning process become very challenging (Li et al., 2021). Loss function for chaotic and multistable dynamics reconstruction. To effectively train RNNs and evaluate reconstruction, it is crucial to carefully choose a proper loss function. The Mean Squared Error (MSE) is a commonly used loss function for reconstruction tasks. MSE is derived under the assumption of Gaussian noise, and it may not be the most appropriate choice when dealing with chaotic systems or multistable dynamics, where the underlying noise characteristics may deviate from Gaussian distribution assumptions. It is not suitable as a test loss for chaotic dynamical systems due to their unpredictable behavior and abrupt changes (Wood, 2010). In multistable systems, where there are multiple stable states, MSE may struggle to distinguish between these states. The loss function may not adequately penalize deviations between different attractors, leading to a less accurate reconstruction of the system’s multistable behavior. Despite proposed alternatives, challenges persist, and an optimal loss function for reconstructing chaotic or multistable dynamics is still lacking (Ciampiconi et al., 2023).
@@
Randomly choosing the internal network parameters. The general idea of randomly choosing the internal parameters of neural networks is studied in random feature models (including deep architectures). Barron (1993) and Rahimi and Recht (2008) developed the basic theory, and Gallicchio and Scardapane (2020) provide a review. Reservoir computing (also called echo-state networks, cf. Jaeger and Haas (2004)) is potentially the closest idea to what we are proposing here. In a reservoir computer, the internal weights are also randomly sampled, and a recurrent, time-delayed model is constructed to approximate a given dynamical system. This type of architecture has been used successfully to model chaotic systems (cf. Pathak et al. (2018); Gauthier et al. (2021); Bollt (2021)). While there are similarities to the recurrent architecture (cf. Lukoševičius and Jaeger (2009)), the concept, as well as the architecture of a reservoir computer, is often treated separately from classic recurrent neural networks that are trained with backpropagation-in-time. In our work, we directly compute all parameters of classical recurrent neural networks without the time-delay component present in reservoir computers.
Koopman operator theory. The Koopman operator is an object associated to every dynamical system. It evolves observables of the state of the system in time. This evolution is linear, which is the main reason the operator is employed and studied extensively for modeling dynamical systems (Mezić, 2005, 2013; Korda and Mezić, 2018b; Gallos et al., 2024). Many numerical approximation algorithms exist (Schmid, 2010; Williams et al., 2015a; Li et al., 2017; Mezic, 2020; Schmid, 2022). The dictionary for the approximation of the Koopman operator has been constructed with neural networks using gradient descent (Li et al., 2017) and using random features (Salam et al., 2022). Reservoir computing has also been related to Koopman operator approximation by Bollt (2021); Gulina and Mauroy (2020). To our knowledge, the relation of the Koopman operator to the weight matrices of recurrent neural networks has not been observed before. This is what we discuss in this work. We also provide a data-dependent probability distribution for the hidden parameters, which is the strongest deviation from the data-agnostic distributions (e.g., normal, uniform) typically used in reservoir computing.
@@ 3 Mathematical framework
We start by defining a general framework of RNNs and the underlying dynamical system before introducing sampling and its connection to the Koopman operator. Let $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$ be an input space, $\mathcal{Y}\subseteq\mathbb{R}^{d_{y}}$ an output space, and $\mathcal{H}\subseteq\mathbb{R}^{d_{h}}$ a state space. We assume that these spaces are associated with the measures $\mu_{x}$, $\mu_{y}$, and $\mu_{h}$, respectively. The underlying dynamical system is then defined through the evolution operator $F$, where we may be working in an uncontrolled system ${\bm{h}}_{t}=F({\bm{h}}_{t-1})$ or a controlled system ${\bm{h}}_{t}=F({\bm{h}}_{t-1},{\bm{x}}_{t})$. We also denote the input dataset $X=[{\bm{x}}_{1},{\bm{x}}_{2},\dots,{\bm{x}}_{N}]$ and the dataset of observations $Y=[{\bm{y}}_{1},{\bm{y}}_{2},\dots,{\bm{y}}_{N}]$. In this paper we are interested in RNNs modelling dynamical systems where the state is observable, and we therefore usually assume access to the dataset $H=[{\bm{h}}_{1},\dots,{\bm{h}}_{N}]$ as well as its copy after one time step $H^{\prime}=[{\bm{h}}^{\prime}_{1},\dots,{\bm{h}}^{\prime}_{N}]$, where ${\bm{h}}^{\prime}_{t}=F({\bm{h}}_{t})$, $t\in\left\{1,\dots,N\right\}$.
@@
Let $W_{h}\in\mathbb{R}^{M\times d_{h}}$, $W_{x}\in\mathbb{R}^{\hat{M}\times d_{x}}$, ${\bm{b}}_{h}\in\mathbb{R}^{M},{\bm{b}}_{x}\in\mathbb{R}^{\hat{M}}$, $C_{h}\in\mathbb{R}^{d_{h}\times M}$, $C_{x}\in\mathbb{R}^{d_{h}\times\hat{M}}$, and $V\in\mathbb{R}^{d_{y}\times d_{h}}$. For time step $t$ and ${\bm{h}}_{0}\in\mathcal{H}$, we define an RNN by
@@ 3.1 Sampling RNN
Sampled neural networks are neural networks where the parameters of the hidden layers are sampled from some distribution, and the last linear layer is either sampled or, more typically, solved through a linear solver. Following ideas from Galaris et al. (2022) and Bolager et al. (2023), we sample the weights and biases of the hidden layers of both $F_{\mathcal{H}}$ and $F_{\mathcal{X}}$ by sampling pairs of points from the domain $\mathcal{H}$ and $\mathcal{X}$, and construct the weights and biases from these pairs of points. One then solves a linear (general) regression problem at the end for the last linear layer that maps to the next state. Concretely, let $\mathbb{P}_{\mathcal{H}}$ and $\mathbb{P}_{\mathcal{X}}$ be probability distributions over $\mathcal{H}^{2}$ and $\mathcal{X}^{2}$ respectively. For each neuron in the hidden layer of $F_{\mathcal{H}}$, sample $({\bm{h}}^{(1)},{\bm{h}}^{(2)})\sim\mathbb{P}_{\mathcal{H}}$ and set the weight $w$ and the bias $b$ of said neuron to
@@
This sampling technique adapts the weights and biases to the underlying domain and constructs weights with direction along the data (see Appendix C for an example how this can be used to interpret the resulting network). Empirically, this has shown to be an improvement over using data-agnostic distributions such as the standard Gaussian one.
One can choose arbitrary probability distributions as $\mathbb{P}_{\mathcal{H}}$ and $\mathbb{P}_{\mathcal{X}}$, with uniform distribution being a common choice. For the supervised setting, Bolager et al. (2023) also proposed a sampling distribution whose density captures the steepest gradients of the target function. For this paper, we sample with densities $p_{\mathcal{H}}$ and $p_{\mathcal{X}}$ proportional to
@@ 3.2 Involving the Koopman operator
The reason for the splitting of matrices $C_{h}$ and $C_{x}$ is the following: the hidden layers $\mathcal{F}_{M}$ and $\mathcal{G}_{\hat{M}}$ map their respective input to a higher dimensional space. In a higher dimensional space, the possibly nonlinear evolution described by $F$ becomes more and more linear (Korda and Mezić, 2018b). This evolution is then captured by the linear transformation of $K$ and $B$ before we map down to the state space through $C$. This also allows us to connect $K$ and $B$ to the Koopman theory applied to the dynamical system. Given a suitable functional space $\mathcal{F}$, the Koopman operator $\mathcal{K}\colon\mathcal{F}\rightarrow\mathbb{R}$ is defined as
@@
The Koopman operator captures the evolution of the dynamical systems in the function space $\mathcal{F}$ and not in the state space itself. In most cases, this makes the operator infinite-dimensional, but in return, a linear operator. For an introduction to the Koopman operator and surrounding theory, see Appendix A. The matrices $K$ and $B$ can then be seen as an approximation of the Koopman operator, and the intuition is that choosing $M$ large enough means we can capture the linear evolution before we map down to the state space.
The matrices $W_{h}$ and $W_{x}$ are found by sampling each row i.i.d. from $\mathbb{P}_{h}$ and $\mathbb{P}_{x}$ respectively. To estimate $C,K,$ and $B$, we use extended dynamic mode decomposition (EDMD) — a classical method to find a finite-dimensional approximation of the Koopman operator. We give a very brief description of EDMD here and give a more thorough introduction in Section A.1. In the uncontrolled setting, this method picks out a dictionary ~$\tilde{\mathcal{F}}_{M}=\{\psi_{1},\dots,\psi_{M}\colon\psi_{i}\colon\mathcal{H}\rightarrow\mathbb{R}\in\tilde{\mathcal{F}}\}$ and estimates the Koopman operator $\mathcal{K}$ using the data $H,H^{\prime}$ by minimizing
$K=\operatorname*{arg\,min}_{\tilde{K}\in\mathbb{R}^{M\times M}}\sum_{n=1}^{N}\lVert\tilde{\mathcal{F}}_{M}({\bm{h}}^{\prime}_{n})-\tilde{K}\tilde{\mathcal{F}}_{M}({\bm{h}}_{n})\rVert,\quad{\bm{h}}^{\prime}_{n}\in H^{\prime},{\bm{h}}_{n}\in H,$
where ~$\tilde{\mathcal{F}}_{M}({\bm{h}})=[\phi_{1}({\bm{h}}),\dots,\phi_{M}({\bm{h}})]^{\mkern-1.5mu\mathsf{T}}$. Letting ~$\tilde{\mathcal{F}}_{M}(H)=[\tilde{\mathcal{F}}_{M}({\bm{h}}_{1}),\dots,\tilde{\mathcal{F}}_{M}({\bm{h}}_{N})]\in\mathbb{R}^{M\times N}$ and ~$\tilde{\mathcal{F}}_{M}(H^{\prime})=[\tilde{\mathcal{F}}_{M}({\bm{h}}^{\prime}_{1}),\dots,\tilde{\mathcal{F}}_{M}({\bm{h}}^{\prime}_{N})]\in\mathbb{R}^{M\times N}$, the approximation can then be written as
$K=\tilde{\mathcal{F}}_{M}(H^{\prime})\,\tilde{\mathcal{F}}_{M}(H)^{+},$ (7)
@@
$[K,B]=\tilde{\mathcal{F}}_{M}(H^{\prime})\,[\tilde{\mathcal{F}}_{M}(H),\tilde{\mathcal{G}}_{\hat{M}}(X)]^{+},$
where $\tilde{\mathcal{G}}_{\hat{M}}$ is the second dictionary mapping from $\mathbb{R}^{d_{x}}$ to $\mathbb{R}^{\hat{M}}$. For more on the Koopman operator in the controlled setting, as well as EDMD, see Section A.2. Regardless of whether uncontrolled or controlled, the mapping $C$ projects down to the state space from the high dimensional dictionary space and is approximated by minimizing $\lVert H-C\tilde{\mathcal{F}}_{M}(H)\rVert$, hence
$C=H\tilde{\mathcal{F}}_{M}(H)^{+}.$
We connect the EDMD algorithm to our recurrent network in Equation 6 by setting the functions $\tilde{\mathcal{F}}$ and $\tilde{\mathcal{G}}$ as hidden layers, i.e., ~$\tilde{\mathcal{F}}_{M}({\bm{h}})=\mathcal{F}_{M}({\bm{h}})=\sigma(W_{h}\,{\bm{h}}+{\bm{b}}_{h})$ and $\tilde{\mathcal{G}}_{\hat{M}}({\bm{x}})=\mathcal{G}_{\hat{M}}({\bm{x}})=\sigma(W_{x}\,{\bm{x}}+{\bm{b}}_{x})$. With this we see that approximating $K$ and $B$ in Equation 6 can be seen as a Koopman approximation with the dictionary being a hidden layer with $M$ and $\hat{M}$ neurons respectively. This connection highlights the benefit of operating in a higher dimensional space. It also allows us to make use of Koopman theory in the next section as the EDMD approximation is known to converge to $\mathcal{K}$. Finally, it is important to notice that the resulting function in Equation 6 consists of two neural networks applied to the previous state and input. The hidden layers are sampled, and the outer matrices are constructed using linear solvers. Hence, Equation 6 is a sampled recurrent neural network. The methods above can be summarized by sampling weights in Algorithm 1, constructing the RNN in Algorithm 2, prediction for uncontrolled systems in Algorithm 3, and model predictive control in Algorithm 4.

### v2

@@ 1 Introduction
Recurrent neural networks (RNNs) are notoriously difficult to train because their loss gradients backpropagated in time tend to saturate or diverge during training, commonly referred to as the Exploding and Vanishing Gradient Problem (EVGP) (Pascanu et al., 2013; Schmidt et al., 2019). To alleviate these problems and improve the computational load of training such networks, we consider a model that completely avoids iterative gradient-descent optimization. We propose to sample the hidden layer parameters of the RNN at random, before solving for the outer linear layer by least-squares methods. Our first major contribution is that we consider data-dependent distributions for the weights and biases, which improves upon the accuracy and interpretability compared to data-agnostic distributions. Our second major contribution adds more structure to the outer layer and improve the models performance even further. We achieve this by connecting RNNs and the Koopman operator (cf. Korda and Mezić (2018a)) through a linear state-space model in a higher-dimensional space, i.e.,
@@
Here the matrix $K$ maps states from ${\bm{z}}_{t-1}$ to ${\bm{z}}_{t}$, with the input ${\bm{x}}_{t}$ affecting this dynamic through the matrix $B$, and the observations are related to the states through the linear map $V$. By using a non-linear neural network to map the low-dimensional state ${\bm{h}}$ into a higher dimension, through ${\mathcal{F}}({\bm{h}}):=\sigma(W{\bm{h}}+{\bm{b}})$, and, similarly, to map the input ${\bm{x}}$ to a high-dimensional input ${\mathcal{G}}({\bm{x}})=\sigma_{x}(W_{x}{\bm{x}}+{\bm{b}}_{x})$, turns the linear SSM into a non-linear recurrent neural network. The final model is illustrated in Figure 1. This connection gives structure and interpretability to the outer layer of RNNs, as well as tools to analyze the RNNs stability through the spectral properties of the Koopman operator.
@@
The combination of sampling hidden layers and applying Koopman theory alleviates all issues related to the backpropagation of gradients and connects the idea of recurrent architectures to the Koopman operator of the underlying dynamical system. The latter allows us to prove convergence results about the approximation quality w.r.t. the number of neurons in ${\mathcal{F}}$. We demonstrate the performance of this approach on several challenging examples with synthetic and real data, and show comparable results to networks trained with backpropagation. Finally, we touch upon the models current challenges such as applying it to high dimensional data.
@@ 2 Related work
Exploding and vanishing gradients. For all major types of RNNs, including LSTMs and GRUs, the dynamics and loss gradients of RNNs are closely linked. If the RNN dynamics converge to a stable fixed point or cycle, loss gradients will remain bounded, but they may vanish (Mikhaeil et al., 2022). Yet, established remedies (Hochreiter and Schmidhuber, 1997; Schmidt et al., 2019) can be used to effectively prevent their gradients from vanishing. However, in chaotic dynamics, gradients invariably explode, posing a challenge that cannot be mitigated through RNN architectural adjustments, regularization, or constraints; instead, it necessitates addressing the problem during the training process (Mikhaeil et al., 2022). Bifurcations may also contribute to sudden jumps in loss observed during RNN training, potentially hindering the training process severely (Doya et al., 1992; Eisenmann et al., 2023). In Eisenmann et al. (2023), it has been demonstrated that specific bifurcations in ReLU-based RNNs are always associated with EVGP during training. Therefore, to harness the full potential of RNNs, the training algorithm needs careful design to tackle challenges posed by bifurcations and the possible emergence of EVGP.
Curse of memory. The existence of long-term memory adversely affects the learning process of RNNs (Bengio et al., 1994; Hochreiter et al., 2001; Li et al., 2021). This negative impact is captured by the concept of the “curse of memory”, which states that when long-term memory is present in the data, approximating relationships demands an exponentially large number of neurons, resulting in a significant slowdown in learning dynamics. Specifically, when the target relationship includes long-term memory, both the approximation and optimization of the learning process become very challenging (Li et al., 2021).
Loss function for chaotic and multistable dynamics reconstruction. To effectively train RNNs and evaluate reconstruction, it is crucial to carefully choose a proper loss function. The Mean Squared Error (MSE) is a commonly used loss function for reconstruction tasks. MSE is derived under the assumption of Gaussian noise, and it may not be the most appropriate choice when dealing with chaotic systems or multistable dynamics, where the underlying noise characteristics may deviate from Gaussian distribution assumptions. It is not suitable as a test loss for chaotic dynamical systems due to their unpredictable behavior and abrupt changes (Wood, 2010). In multistable systems, where there are multiple stable states, MSE may struggle to distinguish between these states. The loss function may not adequately penalize deviations between different attractors, leading to a less accurate reconstruction of the system’s multistable behavior. Despite proposed alternatives, challenges persist, and an optimal loss function for reconstructing chaotic or multistable dynamics is still lacking (Ciampiconi et al., 2023).
@@
Randomly choosing the internal network parameters. The general idea of randomly choosing the internal parameters of neural networks is studied in random feature models (including deep architectures).Barron (1993); Rahimi and Recht (2008) developed the basic theory, and Gallicchio and Scardapane (2020) provide a review. Reservoir computing (also called echo-state networks, cf. Jaeger and Haas (2004)) is potentially the closest idea to what we are proposing here. In a reservoir computer, the internal weights are also randomly sampled, and a recurrent, time-delayed model is constructed to approximate a given dynamical system. This type of architecture has been used successfully to model chaotic systems (cf. Pathak et al. (2018); Gauthier et al. (2021)). While there are similarities to the recurrent architecture (cf. Lukoševičius and Jaeger (2009)), the concept, as well as the architecture of a reservoir computer, is often treated separately from classic recurrent neural networks that are trained with backpropagation-in-time. In our work, we directly compute all parameters of classical recurrent neural networks without the time-delay component present in reservoir computers.
Koopman operator theory. The Koopman operator is an object associated to every dynamical system. It evolves observables of the state of the system in time. This evolution is linear, which is the main reason the operator is employed and studied extensively for modeling dynamical systems (Mezić, 2005, 2013; Korda and Mezić, 2018b). Many numerical approximation algorithms exist (Schmid, 2010; Williams et al., 2015a; Li et al., 2017; Mezic, 2020; Schmid, 2022). The dictionary for the approximation of the Koopman operator has been constructed with neural networks using gradient descent (Li et al., 2017) and using random features (Salam et al., 2022). Reservoir computing has also been related to Koopman operator approximation by Bollt (2021); Gulina and Mauroy (2020). To our knowledge, the relation of the Koopman operator to the weight matrices of recurrent neural networks has not been observed before. This is what we discuss in this work. We also provide a data-dependent probability distribution for the hidden parameters, which is the strongest deviation from the data-agnostic distributions (e.g., normal, uniform) typically used in reservoir computing.
@@ 3 Mathematical framework
We start by defining a general framework of recurrent neural networks (RNNs) and the underlying dynamical system before introducing sampling and its connection to the Koopman operator. Let $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$ be an input space, $\mathcal{Y}\subseteq\mathbb{R}^{d_{y}}$ an output space, and $\mathcal{H}\subseteq\mathbb{R}^{d_{h}}$ a state space. We assume that these spaces are associated with the measures $\mu_{x}$, $\mu_{y}$, and $\mu_{h}$, respectively. The underlying dynamical system is then defined through the evolution operator $F$, where we may be working in an uncontrolled system ${\bm{h}}_{t}=F({\bm{h}}_{t-1})$ or a controlled system ${\bm{h}}_{t}=F({\bm{h}}_{t-1},{\bm{x}}_{t})$. We also denote the input dataset $X=[{\bm{x}}_{1},{\bm{x}}_{2},\dots,{\bm{x}}_{N}]$ and the dataset of observations $Y=[{\bm{y}}_{1},{\bm{y}}_{2},\dots,{\bm{y}}_{N}]$. In this paper we are interested in recurrent neural networks modelling dynamical systems where the state is observable, and we therefore usually assume access to the dataset $H=[{\bm{h}}_{1},\dots,{\bm{h}}_{N}]$ as well as its copy after one time step $H^{\prime}=[{\bm{h}}^{\prime}_{1},\dots,{\bm{h}}^{\prime}_{N}]$, where ${\bm{h}}^{\prime}_{n}=F({\bm{h}}_{n})$, $n\in\left\{1,\dots,N\right\}$.
@@
Let $W_{h}\in\mathbb{R}^{M\times d_{h}}$, $W_{x}\in\mathbb{R}^{\hat{M}\times d_{x}}$, ${\bm{b}}_{h}\in\mathbb{R}^{M},{\bm{b}}_{x}\in\mathbb{R}^{\hat{M}}$, $C_{h}\in\mathbb{R}^{d_{h}\times M}$, $C_{x}\in\mathbb{R}^{d_{h}\times\hat{M}}$, and $V\in\mathbb{R}^{d_{y}\times d_{h}}$. For time step $t$ and ${\bm{h}}_{0}\in\mathcal{H}$, we define a recurrent neural network (RNN) by
@@ 3.1 Sampling RNN
Sampled neural networks are neural networks where the parameters of the hidden layers are sampled from some distribution, and the last linear layer is either sampled or, more typically, solved through a linear solver. Following Bolager et al. (2023), we sample the weights and biases of the hidden layers of both $F_{\mathcal{H}}$ and $F_{\mathcal{X}}$ by sampling pairs of points from the domain $\mathcal{H}$ and $\mathcal{X}$, and construct the weights and biases from these pairs of points. One then solves a linear (general) regression problem at the end for the last linear layer that maps to the next state. Concretely, let $\mathbb{P}_{\mathcal{H}}$ and $\mathbb{P}_{\mathcal{X}}$ be probability distributions over $\mathcal{H}^{2}$ and $\mathcal{X}^{2}$ respectively. For each neuron in the hidden layer of $F_{\mathcal{H}}$, sample $({\bm{h}}^{(1)},{\bm{h}}^{(2)})\sim\mathbb{P}_{\mathcal{H}}$ and set the weight $w$ and the bias $b$ of said neuron to
@@
This sampling technique adapts the weights and biases to the underlying domain and constructs weights with direction along the data (see Appendix C for an example how this can be used to interpret the resulting network). Empirically, this has shown to be an improvement over using data-agnostic distributions such as the standard Gaussian one. One can choose arbitrary probability distributions as $\mathbb{P}_{\mathcal{H}}$ and $\mathbb{P}_{\mathcal{X}}$, with uniform distribution being a common choice. For the supervised setting, Bolager et al. (2023) also proposed a sampling distribution whose density captures the steepest gradients of the target function. For this paper, we sample with densities $p_{\mathcal{H}}$ and $p_{\mathcal{X}}$ proportional to
@@ 3.2 Involving the Koopman operator
The reason for the splitting of matrices $C_{h}$ and $C_{x}$ is the following: the hidden layers $\mathcal{F}_{M}$ and $\mathcal{G}_{\hat{M}}$ map their respective input to a higher dimensional space. In a higher dimensional space, the possibly nonlinear evolution described by $F$ becomes more and more linear (Korda and Mezić, 2018b). This evolution is then captured by $K$ and $B$ before we map down to the state space through $C$. This also allows us to connect $K$ and $B$ to the Koopman theory applied to the dynamical system. Given a suitable functional space $\mathcal{F}$, the Koopman operator $\mathcal{K}\colon\mathcal{F}\rightarrow\mathbb{R}$ is defined as
@@
The Koopman operator captures the evolution of the dynamical systems in the function space $\mathcal{F}$ and not in the state space itself. In most cases, this makes the operator infinite-dimensional, but in return, it is a linear operator. For an introduction to the Koopman operator and surrounding theory, see Appendix A. The matrices $K$ and $B$ can then be seen as an approximation of the Koopman operator, and the intuition is that choosing $M$ large enough means we can capture the linear evolution before we map down to the state space.
The matrices $W_{h}$ and $W_{x}$ are found by sampling each row i.i.d. from $\mathbb{P}_{h}$ and $\mathbb{P}_{x}$ respectively. To estimate $C,K,$ and $B$, we use extended dynamic mode decomposition (EDMD) — a classical method to find a finite-dimensional approximation of the Koopman operator. We give a very brief description of EDMD here and give a more thorough introduction in Section A.1. In the uncontrolled setting, this method picks out a dictionary $\mathcal{F}_{M}=\{\psi_{1},\dots,\psi_{M}\colon\psi_{i}\colon\mathcal{H}\rightarrow\mathbb{R}\in\mathcal{F}\}$ and estimates the Koopman operator $\mathcal{K}$ using the data $H,H^{\prime}$ by minimizing
$K=\operatorname*{arg\,min}_{\tilde{K}\in\mathbb{R}^{M\times M}}\sum_{n=1}^{N}\lVert\mathcal{F}_{M}({\bm{h}}^{\prime}_{n})-\tilde{K}\mathcal{F}_{M}({\bm{h}}_{n})\rVert,\quad{\bm{h}}^{\prime}_{n}\in H^{\prime},{\bm{h}}_{n}\in H,$
where $\mathcal{F}_{M}({\bm{h}})=[\phi_{1}({\bm{h}}),\dots,\phi_{M}({\bm{h}})]^{\mkern-1.5mu\mathsf{T}}$. Letting $\mathcal{F}_{M}(H)=[\mathcal{F}_{M}({\bm{h}}_{1}),\dots,\mathcal{F}_{M}({\bm{h}}_{N})]\in\mathbb{R}^{M\times N}$ and $\mathcal{F}_{M}(H^{\prime})=[\mathcal{F}_{M}({\bm{h}}^{\prime}_{1}),\dots,\mathcal{F}_{M}({\bm{h}}^{\prime}_{N})]\in\mathbb{R}^{M\times N}$, the approximation can then be written as
$K=\mathcal{F}_{M}(H^{\prime})\,\mathcal{F}_{M}(H)^{+},$ (7)
@@
$[K,B]=\mathcal{F}_{M}(H^{\prime})\,[\mathcal{F}_{M}(H),\mathcal{G}_{\hat{M}}(X)]^{+},$
where $\mathcal{G}_{\hat{M}}$ is the second dictionary mapping from $\mathbb{R}^{d_{x}}$ to $\mathbb{R}^{\hat{M}}$. For more on the Koopman operator in the controlled setting, as well as EDMD, see Section A.2. Regardless of whether uncontrolled or controlled, the mapping $C$ projects down to the state space from the high dimensional dictionary space and is approximated by minimizing $\lVert H-C\mathcal{F}_{M}(H)\rVert$, hence
$C=H\mathcal{F}_{M}(H)^{+}.$
We connect the EDMD algorithm to our recurrent network in Equation 6 by choosing $\mathcal{F}_{M}({\bm{h}})=\sigma(W_{h}\,{\bm{h}}+{\bm{b}}_{h})$ and $\mathcal{G}_{\hat{M}}({\bm{x}})=\sigma(W_{x}\,{\bm{x}}+{\bm{b}}_{x})$. With this setting, we see that approximating $K$ and $B$ in Equation 6 can be seen as a Koopman approximation with the dictionary being a hidden layer with $M$ and $\hat{M}$ neurons respectively. This connection highlights the benefit of operating in a higher dimensional space. It also allows us to make use of Koopman theory in the next section as the EDMD approximation is known to converge to $\mathcal{K}$. Finally, it is important to notice that the resulting function in Equation 6 consists of two neural networks applied to the previous state and input. The hidden layers are sampled, and the outer matrices are constructed using linear solvers. Hence, Equation 6 is a sampled recurrent neural network. The methods above can be summarized by sampling weights in Algorithm 1, constructing the RNN in Algorithm 2, prediction for uncontrolled systems in Algorithm 3, and model predictive control in Algorithm 4.